8. Assume that the number of student complaints that arrive at deans office can be modeled as a Poisson random variable. Also assume that on the average there are 5 calls per hour. a) What is the probability that there are exactly 8 complaints in one hour? b) What is the probability that there are 3 or fewer complaints in one hour? c) What is the probability that there are exactly 12 complaints in two hours? d) What is the probability that there are exactly 4 complaints in 30 minutes?
Answer #1

SOLUTION

: X = 5 per hour.

formula: P)

5 * 5, 0.0653 є -=

---------------------------------------------------------------------------------------

{color{Blue} mathbf{b)}};P(Xleq 3)=P(X=0)+P(X=1)+P(X=2)+P(X=3)

1! 2! 3! 0!

5 25 125

P(X < 3) 0.2650

---------------------------------------------------------------------------------------

c) X = 2 * 5 = 10 per tuo hour.

P(X-12) = e-10 * 1012 121 = 0.0948

---------------------------------------------------------------------------------------

d) λ--* 5-2.5 per 30 minutes

P (X = 4) = е-2.5 * 2.54 0.1336 4!